# Provable Ordering and Continuity in Vision-Language Pretraining for Generalizable Embodied Agents

### v1

@@ Abstract
Pre-training vision-language representations on human action videos has emerged as a promising approach to reduce reliance on large-scale expert demonstrations for training embodied agents. However, prior methods often employ time contrastive learning based on goal-reaching heuristics, progressively aligning language instructions from the initial to the final frame. This overemphasis on future frames can result in erroneous vision-language associations, as actions may terminate early or include irrelevant moments in the end. To address this issue, we propose Action Temporal Coherence Learning (AcTOL) to learn ordered and continuous vision-language representations without rigid goal-based constraint. AcTOL treats a video as a continuous trajectory where it (1) contrasts semantic differences between frames to reflect their natural ordering, and (2) imposes a local Brownian bridge constraint to ensure smooth transitions across intermediate frames. Extensive imitation learning experiments across varying numbers of demonstrations show that the pretrained features significantly enhance downstream manipulation tasks by up to 49% with high robustness to different linguistic styles of instructions, offering a viable pathway toward generalized embodied agents. Our code is available here.
@@ 1 Introduction
The long-term vision for embodied intelligence (Mu et al., 2023; Liu et al., 2024) is to create systems that seamlessly perceive and interact with the world around them. Achieving this requires agents that integrate vision and language to understand their surroundings, interpret human instructions, and autonomously plan actions for complex tasks. Current end-to-end approaches achieve policy learning through direct vision-language-action mapping  (Zitkovich et al., 2023; Cheang et al., 2024; Kim et al., 2024). However, the inherent unpredictability of physical environments, including unseen scenarios and dynamic object interactions, constrains these solutions by requiring massive, high-quality robotic trajectories with action annotations, which are costly to collect. To mitigate this, recent research has leveraged large-scale, readily available egocentric human action videos  (Goyal et al., 2017; Damen et al., 2018; Grauman et al., 2022) for pre-training. Although these out-of-domain videos often lack low-level action details and contain noise, their diverse human-object interactions and task instructions provide valuable prior knowledge. This enables the pre-trained representations to be more effectively transferred to novel tasks with fewer demonstrations, reducing reliance on large-scale robotic datasets while preserving strong generalization capabilities.
A promising approach for vision-language pre-training from human action videos leverages the concept of time contrastive learning  (Sermanet et al., 2018) to capture temporally consistent visual representations. In this framework, language serves as the guiding goal, with semantic alignment between the language and chronologically later frames in the video  (Nair et al., 2022; Ma et al., 2023a; Li et al., 2024). However, this semantic alignment approach relies on a rigid assumption that action videos adhere to a specific principle: actions progressively approach the target instruction from the initial frame to the final one. Such assumption can be easily violated in real-world human action videos, which are typically annotated at a coarse-grained level and riddled with noise. The start and end points of actions are often ambiguous, and the progression may not consistently move toward the goal but instead exhibit fluctuations and detours. As a result, these methods struggle with misleading semantic alignment, leading to inaccurate vision-language relationships.
@@
Figure 1: Comparison of existing goal-reaching pre-training strategies and the proposed AcTOL approach. The learned multi-modal representations can be effectively transferred to downstream language-conditioned robot manipulation tasks, exhibiting robustness to diverse instruction and linguistic variations.
@@
To address the aforementioned issues, we propose Action Temporal Coherence Learning (AcTOL), a novel approach designed to implicitly capture the ordering and continuity of video actions without relying on rigid assumptions, while providing strong theoretical guarantees. Unlike previous approaches that focus on goal-directed semantic alignment, AcTOL introduces a Vision-Language Ordering (VLO) loss. This loss leverages the intrinsic temporal coherence of videos, contrasting frames against each other based on their relative temporal distance, theoretically ensuring that the semantic alignment between frames reflects their temporal ordering and continuity throughout the entire sequence. However, the VLO loss does not explicitly enforce the continuity of the visual features themselves, and under conditions with variations in frame content and noise, it can lead to suboptimal local consistency of the visual features. To address this, AcTOL introduces a Brownian bridge constraint over the video, treating video frames as a Brownian bridge process. This approach imposes a structured, continuous flow on the visual representations, ensuring that the model learns more consistent and stable intermediate states, further enhancing the continuity of the visual representations and improving the stability of their alignment with language instruction. Further theoretical analysis suggests that these properties also contribute to the model’s resilience to language perturbations, a crucial trait for real-world applications. To validate the generalization ability of AcTOL on embodied agents, we conducted extensive language-conditioned imitation learning experiments in two simulation environments. The results show that AcTOL outperforms previous methods up to $49.0\%$ with $5$ (limited number of) expert demonstrations. Additionally, we performed an analysis of language-conditioned visual rewards on several real-world action videos. The findings reveal that the ordering and continuity of AcTOL enable it to serve as a promising reward function, generating dense rewards that align well with the given instructions.
2 Related Work
Given the success of large-scale pre-training in the vision and language research communities  (Brown et al., 2020; Liu et al., 2023), many studies have attempted to extend this paradigm to the field of robotics. Some work leverage massive robotic trajectory data  (Collaboration et al., 2023) for pre-training, aiming to establish a unified mapping from the perception space to the action space  (Zitkovich et al., 2023; Cheang et al., 2024). However, collecting large amounts of high-quality robot trajectory data is extremely costly and time-consuming. Consequently, many studies have begun to explore the use of large-scale, readily available, out-of-domain human action video data to learn generalizable representations that can be transferred to robotic tasks  (Sermanet et al., 2018; Ma et al., 2023b; Radosavovic et al., 2022; Nair et al., 2022; Karamcheti et al., 2023; Ma et al., 2023a; Majumdar et al., 2023; Ye et al., 2024; Zeng et al., 2024; Li et al., 2024). Among these, TCN  (Sermanet et al., 2018), VIP  (Ma et al., 2023b), MVP  (Radosavovic et al., 2022), and VC-1  (Majumdar et al., 2023) focus solely on studying unimodal visual representations, limiting their performance when understanding language instructions is required. R3M  (Nair et al., 2022) employs language and reward models to shape progressive visual representations, while Voltron  (Karamcheti et al., 2023) and MPI  (Zeng et al., 2024) model the transition from the current state to the goal state conditioned on language. However, during training, these approaches freeze the language encoder, using it only to aid in the training of visual representations. As a result, they do not effectively achieve multi-modal representation learning.
Recently, LIV  (Ma et al., 2023a) and DecisionNCE  (Li et al., 2024) have attempted to leverage CLIP  (Radford et al., 2021), a state-of-the-art vision-language model, to train embodied multi-modal representations. LIV treats language instructions as the goals of video actions and aligns the final frame of a video with the corresponding language description. DecisionNCE, on the other hand, views language as the transition from the initial state to the final state, aligning the difference between the representations of the first and last frames with the language. Their methods rely on goal-directed semantic alignment, which tends to produce suboptimal results under the noise present in real-world videos. In contrast, our approach avoids rigid assumptions, theoretically ensuring that semantic alignment follows the intrinsic temporal continuity and ordering of the video, resulting in more robust and generalizable vision-language representations.
@@ 3 Preliminaries
Language-Conditioned Imitation Learning (LC-IL). The task of LC-IL aims to train an agent to mimic expert behaviors from a given demonstration set $\mathcal{D}_{d}=\{(\mathbf{\tau}_{i},l_{i})\}_{i=1}^{N}$, where $l_{i}\in\mathcal{L}$ represents a task-specific language instruction. Each trajectory $\mathbf{\tau}_{i}\in\mathcal{T}$ consists of a sequence of state-action pairs $\mathbf{\tau}_{i}=\{(\mathbf{s}_{j},\mathbf{a}_{j})\}_{j=1}^{T}$ of the horizon length $T$. In robot manipulation tasks, action $\mathbf{a}_{j}\in\mathcal{A}$ corresponds to the control commands executed by the agent and state $\mathbf{s}_{j}=[\mathbf{p}_{j};\mathbf{v}_{j}]\in\mathcal{S}$ records proprioceptive data $\mathbf{p}_{j}$ (e.g., joint positions, velocities) and visual inputs $\mathbf{o}_{j}\in\mathcal{O}$ (e.g., camera images) at the time step $j$. The objective of LC-IL is to find an optimal language-conditioned policy $\pi^{*}(\mathbf{a}|\mathbf{s},l):\mathcal{S}\times\mathcal{L}\mapsto\mathcal{A}$ via solving the supervised optimization as follows,
$\pi^{*}\in\arg\min_{\pi}\mathbb{E}_{(\tau_{i},l_{i})\sim\mathcal{T}}\left[\frac{1}{T}\sum_{(\mathbf{s}_{j},\mathbf{a}_{j})\sim\tau_{i}}\ell(\pi(\hat{\mathbf{a}}_{j},\mathbf{s}_{j}|l_{i}),\mathbf{a}_{j})\right],$
@@
Table 1: Comparison of different component designs in time contrast learning across mainstream vision-language pre-training.
$\operatorname{Method}$ $\mathcal{P}(\mathcal{O}_{i})$ $\mathcal{N}(\mathcal{O}_{i})$ $\mathfrak{R}(\mathbf{v},\mathbf{l}_{i})$
[…]
$\operatorname{LIV}$ $(o_{T})$ $(o_{T}^{\notin O_{i}})$ $\operatorname{cos}(\mathbf{v},\mathbf{l}_{i})$
$\operatorname{DecisionNCE}$ $(o_{i},o_{j>i})$ $(o_{i}^{\notin O_{i}},o_{j>i}^{\notin O_{i}})$ $\operatorname{cos}(\mathbf{v}_{j}-\mathbf{v}_{i},\mathbf{l}_{i})$
$\operatorname{AcTOL}$ $(o_{i},o_{j\in[T]\setminus\{i\}})$ $(o_{i},o_{k}:d_{i,k}>d_{i,j})$ $-\|\operatorname{cos}(\mathbf{v}_{i},\mathbf{l}_{i})-\operatorname{cos}(\mathbf{v}_{j},\mathbf{l}_{i})\|_{2}$
@@
Address such scalability issues can be achieved by leveraging large-scale, easily accessible human action video datasets $\mathcal{D}_{p}=\{(\mathcal{O}_{i},l_{i})\}_{i=1}^{M}$ (Damen et al., 2018; Grauman et al., 2022), where $\mathcal{O}_{i}=\{o_{j}\}_{j=1}^{T}$ represents a video clip with $T$ frames and $l_{i}$ the corresponding description. Pretraining on such datasets enables policies to rapidly learn visual-language correspondences with minimal expert demonstrations. Mainstream pretraining methods employ time contrastive learning (Sermanet et al., 2018) to fine-tune a visual encoder $\mathcal{\phi}$ and a text encoder $\mathcal{\varphi}$, which project frames and descriptions into a shared $d$-dimensional embedding space, i.e., $\mathbf{v}_{j}=\phi(o_{j})\in\mathbb{R}^{d}$ and $\mathbf{l}_{i}=\varphi(l_{i})\in\mathbb{R}^{d}$. To provide a unified perspective on various pretraining approaches, we formulate them within the objective $\mathcal{L}_{\operatorname{tNCE}}(\phi,\varphi)$:
@@
$\small\mathcal{L}_{\operatorname{tNCE}}$ $=-\mathbb{E}_{\begin{subarray}{c}\scriptstyle o^{+}\sim{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{P}(\mathcal{O}_{i})}\end{subarray}}\log\frac{\exp(\mathfrak{R}(\mathbf{v}^{+},\mathbf{l}_{i}))}{\mathbb{E}_{\scriptstyle o^{-}\sim{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{N}(\mathcal{O}_{i})}}\exp(\mathfrak{R}(\mathbf{v}^{-},\mathbf{l}_{i}))},$
@@
where $\mathbf{v}^{+/-}=\phi(o^{+/-})$. Different pretraining strategies differ in their selection of (1) the positive frame set $\mathcal{P}(\mathcal{O}_{i})$, (2) negative frame set $\mathcal{N}(\mathcal{O}_{i})$; and (3) the semantic alignment scoring function $\mathfrak{R}(\mathbf{v},\mathbf{l}_{i})$ measuring the gap of VL similarities as detailed in Table 1.
@@
Discussion. As motivated by goal-conditioned RL (Andrychowicz et al., 2017), current approaches explicitly select future frames (e.g., DecisionNCE) or the last frame (e.g., LIV) as the goal within the positive set, enforcing their visual embedding to align with the semantics. Likewise, the scoring functions $\mathfrak{R}$ are often designed to maximize this transition direction. However, the pretraining action videos are noisy as actions may terminate early or include irrelevant subsequent actions, which may mislead the encoders and result in inaccurate vision-language association. As detecting precise action boundaries is non-trivial, we argue for a more flexible approach that leverages intrinsic characteristics of actions to guide pretraining.
4 Our Approach
@@
We introduce an action temporal coherence learning (AcTOL) to capture two temporal properties of video actions: ordering and continuity. Ordering was ensured in the vision-language ordering loss (Section 4.1), where the semantic difference between frames reflects their temporal distance, with closer frames exhibiting smaller differences than those further apart. Continuity requires smooth visual transitions between adjacent frames, avoiding abrupt changes and high variance. To achieve this, we model sampled frame intervals as a Brownian bridge process (Section 4.2), penalizing deviations from the expected trajectories. Different from prior works that relies on setting explicit goal frames, the proposed approach implictly explore the global and local structure of actions without imposing rigid constraints.
@@ 4.1 Visual-Language Ordering
To capture the temporal coherence of video actions, we first propose a vision-language ordering (VLO) loss that ensures the semantic alignment between frames reflects their temporal order. Consider an anchor frame $o_{i}\in\mathcal{O}$ with an index $n(i)$ corresponding to its position in the original video. For any given frame pair $(o_{i},o_{j})$, we first define the semantic alignment score $\mathfrak{R}$ to quantify differences in their VL similarities w.r.t a language description $l$ as:
@@
where $\mathbf{v}_{i}=\phi(o_{i})$, $\mathbf{l}=\varphi(l)$. The function $\operatorname{sim}(\cdot,\cdot)$ computes the VL similarity using cosine similarity. To ensure the proposed $\mathfrak{R}$ adhere to the temporal ordering of frames, we construct a negative set $\mathcal{N}_{i,j}$ by selecting $o_{k}\in O$ correspond to frames that are temporally more distant than the positive pair $(o_{i},o_{j})$:
@@
$\small\mathcal{L}_{\operatorname{VLO}}=-\mathbb{E}_{(o_{i},o_{j})\sim O}\log\frac{\exp\left(\mathfrak{R}(\mathbf{v}_{i},\mathbf{v}_{j},\mathbf{l})\right)}{\sum_{o_{k}\in\mathcal{N}_{i,j}}\exp\left(\mathfrak{R}(\mathbf{v}_{i},\mathbf{v}_{k},\mathbf{l})\right)}.$
Notably, our VLO loss does not strictly require $o_{j}$ to be from future timestep for goal-reaching. Instead, we leverage the inherent temporal dynamics in videos, allowing the model to learn the natural ordering in an unsupervised manner with detailed analysis as follows.
Theoretical Analysis. Ordering and sorting properties are well-established in self-supervised learning (Shvetsova et al., 2023; Hu et al., 2021; Zha et al., 2023). Building upon these insights, we formalize the concept of vision-language ordering below.
Definition 4.1 (Vision-Language Ordering).
Let $\{o_{i}\}_{i\in[T]}$ be a sequence of video frames and $l$ the corresponding language description. The representations of the frames are said to satisfy the VLO property for any $10<\delta<$ 1 if $\forall i\in[T]$, and distinct frames $j,k\in[T]\backslash\{i\}$, the following conditions hold:
$\left\{\begin{array}[]{ll}\mathfrak{R}_{i,j,l}>\mathfrak{R}_{i,k,l}+1/{\delta},&\text{if }d_{i,j}<d_{i,k},\\ \left|\mathfrak{R}_{i,j,l}-\mathfrak{R}_{i,k,l}\right|<\delta,&\text{if }d_{i,j}=d_{i,k},\\ \mathfrak{R}_{i,j,l}<\mathfrak{R}_{i,k,l}-1/{\delta},&\text{if }d_{i,j}>d_{i,k},\end{array}\right.$
where $\mathfrak{R}_{i,j,l}$ denotes $\mathfrak{R}\left(\mathbf{v}_{i},\mathbf{v}_{j},\mathbf{l}\right)$ and the temporal distance between frames defined as $|n(i)-n(j)|$ as $d_{i,j}$.
To formalize the temporal ordering constraints, we define the unique sorted set of frame distances from frame $i$ as $\{D_{i,1}<D_{i,2}<\cdots<D_{i,M}\}$, where each $D_{i,m},m\in[M_{i}]$ is obtained by sorting the set $\{d_{i,j}\mid j\in[T]\setminus\{i\}\}$. Additionally, we define the count of frames at each distance level as:
$n_{i,m}:=|\{j\mid d_{i,j}=D_{i,m},\,j\in[T]\setminus\{i\}\}|,$ (2)
This bound characterizes the optimal alignment of VL similarities, ensuring that the learned representations preserve the inherent temporal structure within the video sequence, as guaranteed by the following theorem.
Theorem 4.2.
$\mathcal{L}^{*}$ is a tight lower bound of $\mathcal{L}_{\mathrm{VLO}}$, i.e., $\mathcal{L}_{\mathrm{VLO}}\geq\mathcal{L}^{*}$, and for any $\epsilon>0$, there exists feature embeddings such that $\mathcal{L}_{\mathrm{VLO}}<\mathcal{L}^{*}+\epsilon$. Furthermore, for any $10<\delta<$ 1, there exist $\epsilon>0$such that if $\mathcal{L}_{\mathrm{VLO}}<\mathcal{L}^{*}+\epsilon$, the learned representations satisfy the VLO property.
Proof.
Please refer to Appendix A.1. ∎

### v3

@@ Abstract
Pre-training vision-language representations on human action videos has emerged as a promising approach to reduce reliance on large-scale expert demonstrations for training embodied agents. However, prior methods often employ time contrastive learning based on goal-reaching heuristics, progressively aligning language instructions from the initial to the final frame. This overemphasis on future frames can result in erroneous vision-language associations, as actions may terminate early or include irrelevant moments in the end. To address this issue, we propose Action Temporal Coherence Learning (AcTOL) to learn ordered and continuous vision-language representations without rigid goal-based constraint. AcTOL treats a video as a continuous trajectory where it (1) contrasts semantic differences between frames to reflect their natural ordering, and (2) imposes a local Brownian bridge constraint to ensure smooth transitions across intermediate frames. Extensive imitation learning experiments on both simulated and real robots show that the pretrained features significantly enhance downstream manipulation tasks with high robustness to different linguistic styles of instructions, offering a viable pathway toward generalized embodied agents. Our project page is at https://actol-pretrain.github.io/.
Figure 1: Pretraining on Internet human action videos for robot control, where the video-instruction pairs are noisy and often include irrelevant frames. The red vision-language reward curve demonstrates AcTOL learns to correctly align instruction with action, outperforming previous goal-reaching methods in the presence of distracting content.
@@ 1 Introduction
The long-term vision for embodied intelligence [DBLP:conf/nips/MuZHWDJWDQL23, DBLP:journals/corr/abs-2407-06886] is to create systems that seamlessly perceive and interact with the world around them. Achieving this requires agents that integrate vision and language to understand their surroundings, interpret human instructions, and autonomously plan actions for complex tasks. Current end-to-end approaches achieve policy learning through direct vision-language-action mapping [corl23-rt2, octo, corr24-gr2, corr24-openvla, pi-zero]. However, the inherent unpredictability of physical environments, including unseen scenarios and dynamic object interactions, constrains these solutions by requiring massive, high-quality robotic trajectories with action annotations, which are costly to collect. To mitigate this, recent research has leveraged large-scale, readily available egocentric human action videos [DBLP:conf/iccv/GoyalKMMWKHFYMH17, corr18-epickitchen, cvpr22-ego4d] for pre-training. Although these out-of-domain videos often lack low-level action details and contain noise, their diverse human-object interactions and task instructions provide valuable prior knowledge. This enables the pre-trained representations to be more effectively transferred to novel tasks with fewer demonstrations, reducing reliance on large-scale robotic datasets while preserving strong generalization capabilities.
A promising approach for vision-language pre-training from human action videos leverages the concept of time contrastive learning [icra18-tcn] to capture temporally consistent visual representations, where language serves as the guiding goal, with semantic alignment between the language and chronologically later frames in the video [corl22-r3m, icml23-liv, icml24-decisionnce]. However, this goal-reaching semantic alignment approach relies on a rigid assumption that action videos adhere to a specific principle: actions progressively approach the target instruction from the initial frame to the final one. Such assumption can be easily violated in egocentric human action videos, which are typically annotated at a coarse-grained level and riddled with noise. Figure 1 shows an example video-instruction pair, where the end of the video clip does not correspond to the actual end of the action. As a result, existing methods suffer from misleading semantic alignment, which hampers their ability to learn accurate vision-language relationships.
Figure 2: Comparison of existing goal-reaching pre-training strategies and the proposed AcTOL approach. Our learned multi-modal representations can be effectively transferred to downstream language-conditioned robot manipulation tasks, exhibiting robustness to diverse instruction and linguistic variations.
@@
To address the aforementioned issues, as illustrated in Figure 2, we propose Action Temporal Coherence Learning (AcTOL), a novel approach designed to implicitly capture the ordering and continuity of video actions without relying on rigid assumptions, while providing strong theoretical guarantees. Unlike previous approaches that focus on goal-directed semantic alignment, AcTOL introduces a Vision-Language Ordering (VLO) loss. This loss leverages the intrinsic temporal coherence of videos, contrasting frames against each other based on their relative temporal distance, theoretically ensuring that the semantic alignment between frames reflects their temporal ordering and continuity throughout the entire sequence. However, the VLO loss does not explicitly enforce the continuity of the visual features themselves, and under conditions with variations in frame content and noise, it can lead to suboptimal local consistency of the visual features. To address this, AcTOL introduces a Brownian bridge constraint over the video, treating video frames as a Brownian bridge process. This approach imposes a structured, continuous flow on the visual representations, ensuring that the model learns more consistent and stable intermediate states, further enhancing the continuity of the visual representations and improving the stability of their alignment with language instruction. Further theoretical analysis suggests that these properties also contribute to the model’s resilience to language perturbations, a crucial trait for real-world applications. To evaluate the generalization ability of AcTOL on embodied agents, we conducted extensive language-conditioned imitation learning experiments using both the real-world Unitree D1 robotic arm and two simulation environments. The results demonstrate that AcTOL significantly outperforms prior methods with a limited number of expert demonstrations. Additionally, AcTOL can generate language-conditioned visual rewards from real-world robot videos and remains robust to complex linguistic perturbations, highlighting its potential as a generalizable solution for real-world embodied agents.
@@ 2 Preliminaries
Language-Conditioned Imitation Learning (LC-IL). The task of LC-IL aims to train an agent to mimic expert behaviors from a given robot demonstration set $\mathcal{D}_{\text{robot}}=\{(\mathbf{\tau}_{i},l_{i})\}_{i=1}^{N_{r}}$, where $l_{i}\in\mathcal{L}$ represents a task-specific language instruction. Each trajectory $\mathbf{\tau}_{i}\in\mathcal{T}$ consists of a sequence of state-action pairs $\mathbf{\tau}_{i}=\{(\mathbf{s}_{t},\mathbf{a}_{t})\}_{t=1}^{T}$ of the horizon length $T$. In robot manipulation tasks, action $\mathbf{a}_{t}\in\mathcal{A}$ corresponds to the control commands executed by the agent and state $\mathbf{s}_{t}=[\mathbf{p}_{t};\mathbf{o}_{t}]\in\mathcal{S}$ records proprioceptive data $\mathbf{p}_{t}$ (e.g., joint positions, velocities) and visual inputs $\mathbf{o}_{t}$ (e.g., camera images) at the time step $t$. The objective of LC-IL is to find an optimal language-conditioned policy $\pi^{*}(\mathbf{a}|\mathbf{s},l):\mathcal{S}\times\mathcal{L}\mapsto\mathcal{A}$ via solving the supervised optimization as follows,
$\pi^{*}\in\arg\min_{\pi}\mathbb{E}_{(\tau_{i},l_{i})\sim\mathcal{T}}\left[\frac{1}{T}\sum_{(\mathbf{s}_{t},\mathbf{a}_{t})\sim\tau_{i}}\ell(\pi(\hat{\mathbf{a}}_{t},\mathbf{s}_{t}|l_{i}),\mathbf{a}_{t})\right],$
@@
Address such scalability issues can be achieved by leveraging large-scale, easily accessible human action video datasets [corr18-epickitchen, cvpr22-ego4d] $\mathcal{D}_{\text{human}}=\{(\mathcal{O}_{i},l_{i})\}_{i=1}^{N_{h}}$ , where $\mathcal{O}_{i}=\{o_{t}\}_{t=1}^{T}$ represents a video clip with $T$ frames and $l_{i}$ the corresponding description. Pretraining on such datasets enables policies to rapidly learn visual-language correspondences with minimal expert demonstrations. Mainstream pretraining methods employ time contrastive learning [icra18-tcn] to fine-tune a visual encoder $\mathcal{\phi}$ and a text encoder $\mathcal{\varphi}$, which project frames and descriptions into a shared $d$-dimensional embedding space, i.e., $\mathbf{v}_{t}=\phi(o_{t})\in\mathbb{R}^{d}$ and $\mathbf{l}_{i}=\varphi(l_{i})\in\mathbb{R}^{d}$. To provide a unified perspective on various pretraining approaches, we formulate them within the objective $\mathcal{L}_{\operatorname{tNCE}}(\phi,\varphi)$:
@@
$\mathcal{L}_{\operatorname{tNCE}}$ $=-\mathbb{E}_{\begin{subarray}{c}\scriptstyle o^{+}\sim{\color[rgb]{0,0,0}\mathcal{P}(\mathcal{O}_{i})}\end{subarray}}\log\frac{\exp(\mathfrak{R}(\mathbf{v}^{+},\mathbf{l}_{i}))}{\mathbb{E}_{\scriptstyle o^{-}\sim{\color[rgb]{0,0,0}\mathcal{N}(\mathcal{O}_{i})}}\exp(\mathfrak{R}(\mathbf{v}^{-},\mathbf{l}_{i}))},$
@@
where $\mathbf{v}^{+/-}=\phi(o^{+/-})$. Different pretraining strategies differ in their selection of (1) the positive frame set $\mathcal{P}(\mathcal{O}_{i})$, (2) negative frame set $\mathcal{N}(\mathcal{O}_{i})$; and (3) the semantic alignment scoring function $\mathfrak{R}(\mathbf{v},\mathbf{l}_{i})$ measuring the gap of VL similarities.
@@
As motivated by goal-conditioned RL [nips17-her], current approaches explicitly select future frames (e.g., R3M [corl22-r3m], DecisionNCE [icml24-decisionnce]) or the last frame (e.g., LIV [icml23-liv]) as the goal within the positive frame set, enforcing their visual embedding to align with the semantics. Likewise, the scoring functions $\mathfrak{R}$ are often designed to maximize this transition direction. However, the pretraining action videos are noisy as actions may terminate early or include irrelevant subsequent actions, which may mislead the encoders and result in inaccurate vision-language association. As detecting precise action boundaries is non-trivial, we argue for a more flexible approach that leverages intrinsic characteristics of actions to guide pertaining.
3 Our Approach: AcTOL
@@
We introduce an action temporal coherence learning (AcTOL) to capture two temporal properties of video actions: ordering and continuity. Ordering was ensured in the vision-language ordering loss (Section 3.1), where the semantic difference between frames reflects their temporal distance, with closer frames exhibiting smaller differences than those further apart. Continuity requires smooth visual transitions between adjacent frames, avoiding abrupt changes and high variance. To achieve this, we model sampled frame intervals as a Brownian bridge process (Section 3.2), penalizing deviations from the expected trajectories. Different from prior works that relies on setting explicit goal frames, the proposed approach implicitly explore the global and local structure of actions without imposing rigid constraints.
@@ 3.1 Visual-Language Ordering
To capture the temporal coherence of video actions, we first propose a vision-language ordering (VLO) loss that ensures the semantic alignment between frames reflects their temporal order. Since the VLO loss is applied within each video individually, we henceforth write $\mathcal{O}_{i},l_{i}$ as $\mathcal{O},l$ for simplicity. Consider an anchor frame $o_{i}\in\mathcal{O}$ with an index $n(i)$ corresponding to its position in the original video. For any given frame pair $(o_{i},o_{j})$, we first define the semantic alignment score $\mathfrak{R}$ to quantify differences in their VL similarities w.r.t a language description $l$ as:
@@
where $\mathbf{v}_{i}=\phi(o_{i})$, $\mathbf{l}=\varphi(l)$. The function $\operatorname{sim}(\cdot,\cdot)$ computes the VL similarity using cosine similarity. To ensure the proposed $\mathfrak{R}$ adhere to the temporal ordering of frames, we construct a negative set $\mathcal{N}_{i,j}$ by selecting $o_{k}\in\mathcal{O}$ correspond to frames that are temporally more distant from $o_{i}$ than $o_{j}$:
@@
$\mathcal{L}_{\operatorname{VLO}}=-\mathbb{E}_{(o_{i},o_{j})\sim\mathcal{O}}\log\frac{\exp\left(\mathfrak{R}(\mathbf{v}_{i},\mathbf{v}_{j},\mathbf{l})\right)}{\sum_{o_{k}\in\mathcal{N}_{i,j}}\exp\left(\mathfrak{R}(\mathbf{v}_{i},\mathbf{v}_{k},\mathbf{l})\right)}.$
Notably, our VLO loss does not strictly require $o_{j}$ to be from a future timestep for goal-reaching. Instead, we leverage the inherent temporal dynamics in videos, allowing the model to learn the natural ordering in an unsupervised manner.